# Variation of cones of divisors in a family of varieties – Fano type case

### v3

@@ 5.3. Deformation of effective cones
Let $f:X\to T$ be a projective fibration. Suppose that there exists a Zariski dense subset $S\subset T$ such that the fibers $X_{s}$ for $s\in S$ satisfy one of the cases in Theorem 1.2. Then, up to a generic finite base change of $T$, there exists a non-empty open subset $T_{0}\subset T$, such that for any $U\subset T_{0}$, we have
@@
After replacing $T$ by a generic finite base change and possibly shrinking it, we may assume that Theorem 1.6 holds. By Theorem 1.2, after shrinking $T$, we can assume that $X/T$ is of Fano type. If $\mathcal{D}\in{\rm{Eff}}(X/T)$, then we can run a $\mathcal{D}$-MMP$/T$, $X\dashrightarrow X_{n}$, which terminates to $X_{n}$ such that $\mathcal{D}_{X_{n}}$ is nef $/T$. By Theorem 1.6 (1), it is also a $\mathcal{D}_{t}$-MMP on $X_{t}$ for any $t\in T$. This shows that under the identification $N^{1}(X/T)\simeq N^{1}(X_{t})$ (see Theorem 1.6 (3)), we have
@@
Conversely, take $D\in{\rm{Eff}}(X_{t})$. Let $X_{t}\dasharrow Y$ be a $D$-MMP such that $D_{Y}$ is nef on $Y$. Then by Theorem 1.6 (2), it is the restriction of a $\mathcal{D}$-MMP$/T$ on $X$, where $\mathcal{D}$ is a divisor on $X$ such that $[\mathcal{D}_{t}]=[D]$. Hence we have $[\mathcal{D}]\in{\rm{Eff}}(X/T)$. This show the inverse inclusion ${\rm{Eff}}(X/T)\supset{\rm{Eff}}(X_{t})$. ∎
@@ 5.4. Deformation of movable cones and Mori chamber decompositions
If $X$ is of Fano type over $T$ and $[D]\in{\rm{Mov}}(X/T)=\overline{\rm{Mov}}(X/T)$ is a movable divisor, then there exists a $D$-MMP$/T$, $\phi:X\dashrightarrow Y/T$, which terminates at $Y$. Note that $X\dashrightarrow Y$ is isomorphic in codimension $1$ and if $D_{Y}$ is the strict transform of $D$ on $Y$, then $[D_{Y}]\in{\rm{Nef}}(Y/T)$. Conversely, if $\phi:X\dashrightarrow Y$ is isomorphic in codimension $1$, then $\phi^{-1}_{*}{\rm{Nef}}(Y/T)\cap N^{1}(X/T)\subset{\rm{Mov}}(X/T)$. Here the chamber $\phi^{-1}_{*}{\rm{Nef}}(Y/T)\cap N^{1}(X/T)$ consists of $\mathbb{R}$-Cartier divisors which are strict transforms of nef divisors on $Y/T$.
@@
Let $f:X\to T$ be a projective fibration. Suppose that there exists a subset $S\subset T$ such that the fibers $X_{s}$ for $s\in S$ satisfy one of the cases in Theorem 1.2. Then, up to a generic finite base change of $T$, there exists a Zariski open subset $T_{0}\subset T$, such that for any $U\subset T_{0}$, we can identify the Mori chamber decompositions of ${\rm{Mov}}(X_{T_{0}}/T_{0}),{\rm{Mov}}(X_{U}/U)$ and ${\rm{Mov}}(X_{t}),t\in U$ under the natural restriction maps. In particular, we have isomorphisms among movable cones
@@
After replacing $T$ by a generic finite base change and possibly shrinking it, we may assume that Theorem 1.6 holds. By Theorem 1.2, after shrinking $T$, we can assume that $X/T$ is of Fano type. By Theorem 1.6 (1) and (2), an MMP$/T$ on $X$ restricts to an MMP of the same type on each fiber $X_{t}$, and the converse also holds. Therefore, ${\rm{Mov}}(X/T)$ and ${\rm{Mov}}(X_{t})$ shares the same chambers under the identification $N^{1}(X/T)\simeq N^{1}(X_{t})$ (see Theorem 1.6 (3)). ∎
@@ 6. Boundedness of birational models of Fano type varieties
By definition, $\mathcal{S}$ consists of closed fibers of a family of Fano type varieties $X/T$. By Noetherian induction and Theorem 1.4, after replacing $T$ by a generic finite base change and possibly shrinking it, we can assume that there exists a fiber-wise small $\mathbb{Q}$-factorial modification $Y\to X$. Note that ${\rm bcm}(Y_{t})={\rm bcm}(X_{t})$. Replacing $Y$ by $X$, we can assume that $X_{t},t\in T$ are $\mathbb{Q}$-factorial.
After replacing $T$ by a generic finite base change and possibly shrinking it further, we may assume that Theorem 1.6 holds. Note that this replacement does not affect the $\mathbb{Q}$-factoriality of the fibers. By Theorem 1.2 and Noetherian induction, after shrinking $T$, we can assume that $X$ is of Fano type over $T$.
Fix a $X_{t}$ for some $t\in T$, and suppose that $X_{t}\dasharrow Z$ is a birational contraction. As $X_{t}$ is a $\mathbb{Q}$-factorial variety, we have $Z\in{\rm{bc}}(X_{t})_{\rm MMP}$ by Proposition 5.4. By Theorem 1.6 (2), there exists an MMP$/T$, denoted by $X\dasharrow\mathcal{Z}/T$, such that $Z\simeq\mathcal{Z}_{t}$. Note that $\mathcal{Z}/T\in{\rm{bc}}(X/T)$. By Proposition 5.1, ${\rm{bc}}(X/T)$ is a finite set. Thus, ${\rm bcm}(\mathcal{S})$ is bounded by definition. ∎
@@
where the notation follows the same meaning as above. In particular, there exists a divisor $E_{n-1}$ whose discrepancy with respect to $(X,c_{n}D+c_{n}B)$ is 00. By [BCHM10], we can extract $E_{n-1}$ alone to get a variety $\theta_{n}:Y_{n}\to X$. Therefore, we have

### v4

@@ 5.3. Deformation of effective cones
Let $f:X\to T$ be a projective fibration. Suppose that there exists a Zariski dense subset $S\subset T$ such that the fibers $X_{s}$ for $s\in S$ satisfy one of the cases in Theorem 1.2. Then, up to a generically finite base change of $T$, there exists a non-empty open subset $T_{0}\subset T$, such that for any $U\subset T_{0}$, we have
@@
After replacing $T$ by a generically finite base change and possibly shrinking it, we may assume that Theorem 1.6 holds. By Theorem 1.2, after shrinking $T$, we can assume that $X/T$ is of Fano type. If $\mathcal{D}\in{\rm{Eff}}(X/T)$, then we can run a $\mathcal{D}$-MMP/T/T, $X\dashrightarrow X_{n}$, which terminates to $X_{n}$ such that $\mathcal{D}_{X_{n}}$ is nef /T/T. By Theorem 1.6 (2), it is also a $\mathcal{D}_{t}$-MMP on $X_{t}$ for any $t\in T$. This shows that under the identification $N^{1}(X/T)\simeq N^{1}(X_{t})$ (see Theorem 1.6 (4)), we have
@@
Conversely, take $D\in{\rm{Eff}}(X_{t})$. Let $X_{t}\dasharrow Y$ be a $D$-MMP such that $D_{Y}$ is nef on $Y$. Then by Theorem 1.6 (3), it is the restriction of a $\mathcal{D}$-MMP/T/T on $X$, where $\mathcal{D}$ is a divisor on $X$ such that $\mathcal{D}_{t}]=[D]$. Hence we have $\mathcal{D}]\in{\rm{Eff}}(X/T)$. This show the inverse inclusion ${\rm{Eff}}(X/T)\supset{\rm{Eff}}(X_{t})$. ∎
@@ 5.4. Deformation of movable cones and Mori chamber decompositions
If $X$ is of Fano type over $T$ and $D]\in{\rm{Mov}}(X/T)=\overline{\rm{Mov}}(X/T)$ is a movable divisor, then there exists a $D$-MMP/T/T, $\phi:X\dashrightarrow Y/T$, which terminates at $Y$. Note that $X\dashrightarrow Y$ is isomorphic in codimension $1$ and if $D_{Y}$ is the strict transform of $D$ on $Y$, then $D_{Y}]\in{\rm{Nef}}(Y/T)$. Conversely, if $\phi:X\dashrightarrow Y$ is isomorphic in codimension $1$, then $\phi^{-1}_{*}{\rm{Nef}}(Y/T)\cap N^{1}(X/T)\subset{\rm{Mov}}(X/T)$. Here the chamber $\phi^{-1}_{*}{\rm{Nef}}(Y/T)\cap N^{1}(X/T)$ consists of $\mathbb{R}$-Cartier divisors which are strict transforms of nef divisors on $Y/T$.
@@
Let $f:X\to T$ be a projective fibration. Suppose that there exists a subset $S\subset T$ such that the fibers $X_{s}$ for $s\in S$ satisfy one of the cases in Theorem 1.2. Then, up to a generically finite base change of $T$, there exists a Zariski open subset $T_{0}\subset T$, such that for any $U\subset T_{0}$, we can identify the Mori chamber decompositions of ${\rm{Mov}}(X_{T_{0}}/T_{0}),{\rm{Mov}}(X_{U}/U)$ and ${\rm{Mov}}(X_{t}),t\in U$ under the natural restriction maps. In particular, we have isomorphisms among movable cones
@@
After replacing $T$ by a generically finite base change and possibly shrinking it, we may assume that Theorem 1.6 holds. By Theorem 1.2, after shrinking $T$, we can assume that $X/T$ is of Fano type. By Theorem 1.6 (2) and (3), an MMP/T/T on $X$ restricts to an MMP of the same type on each fiber $X_{t}$, and the converse also holds. Therefore, ${\rm{Mov}}(X/T)$ and ${\rm{Mov}}(X_{t})$ shares the same chambers under the identification $N^{1}(X/T)\simeq N^{1}(X_{t})$ (see Theorem 1.6 (4)). ∎
@@ 6. Boundedness of birational models of Fano type varieties
By definition, $\mathcal{S}$ consists of closed fibers of a family of Fano type varieties $X/T$. By Noetherian induction and Theorem 1.4, after replacing $T$ by a generically finite base change and possibly shrinking it, we can assume that there exists a fiber-wise small $\mathbb{Q}$-factorial modification $Y\to X$. Note that ${\rm bcm}(Y_{t})={\rm bcm}(X_{t})$. Replacing $Y$ by $X$, we can assume that $X_{t},t\in T$ are $\mathbb{Q}$-factorial.
After replacing $T$ by a generically finite base change and possibly shrinking it further, we may assume that Theorem 1.6 holds. Note that this replacement does not affect the $\mathbb{Q}$-factoriality of the fibers. By Theorem 1.2 and Noetherian induction, after shrinking $T$, we can assume that $X$ is of Fano type over $T$.
Fix a $X_{t}$ for some $t\in T$, and suppose that $X_{t}\dasharrow Z$ is a birational contraction. As $X_{t}$ is a $\mathbb{Q}$-factorial variety, we have $Z\in{\rm{bc}}(X_{t})_{\rm MMP}$ by Proposition 5.4. By Theorem 1.6 (3), there exists an MMP/T/T, denoted by $X\dasharrow\mathcal{Z}/T$, such that $Z\simeq\mathcal{Z}_{t}$. Note that $\mathcal{Z}/T\in{\rm{bc}}(X/T)$. By Proposition 5.1, ${\rm{bc}}(X/T)$ is a finite set. Thus, ${\rm bcm}(\mathcal{S})$ is bounded by definition. ∎
@@
where the notation follows the same meaning as above. In particular, there exists a divisor $E_{n-1}$ whose discrepancy with respect to $(X,c_{n}D+c_{n}B)$ is 0. By [BCHM10], we can extract $E_{n-1}$ alone to get a variety $\theta_{n}:Y_{n}\to X$. Therefore, we have